OpenStudy (anonymous):

f(x)=3x-7 and g(x)=-2x-6 find (f o g)(4)

OpenStudy (anonymous):

Please help!!!!

OpenStudy (anonymous):

Work from right to left. First, substitute x=4 into the g(x) equation. Then substitute x = g(4) into f(x), g(4) being the value you just found in the first step.

OpenStudy (anonymous):

I'm so confused

OpenStudy (anonymous):

So g(x) = -2x-6 g(4) = -2(4)-6 = -8-6 = -14 f(x) = 3x-7 f(g(4)) = 3(-14)-7 = -42-7 = -49 so (f o g)(4) = -49